Calculate the specific heat capacity of a 3.0kg piece of metal which experiences a temperature rise of 25∘C when it is heated at a rate of 60W for 10minutes, if a total of 3000J heats the surroundings during this process.
1
Expert's answer
2020-03-31T09:38:16-0400
"Pt=mc\Delta T+Q_{lost}"

"(60)(10\cdot 60)=(3)(25)c+3000"

"c=440\frac{J}{kg\degree C}"